OpenStudy (anonymous):

if the base of a prism is rhombus with diagnols 28cm and 96 cm,height of prism is 32 cm then its volume is ?

Directrix (directrix):

@sanjuna The volume of a prism can be found using the formula V=B*h where B is the area of the base of the prism and h is the height of the prism. The Base of this prism is a rhombus. The area of a rhombus is one-half the product of its diagonals. To get the volume of the prism: V = B*h V = (1/2) * (28*96) *32 V = ?